one side of a rectangle is 5 inches and the other side is x inches. find the value of x if the area must be at least 60 square inches. use A=L*W
Answer by oberobic(2304)

We know start with what is given:
A = L*W = 60
...
Pick a side to set = 5.
L = 5
...
So W has to be 60/5 = 12
...
Be careful to answer the asked question, which refers to 'x'.
x = 12
...
Now check your work.
Is one side 5? Yes.
Is the area 60? Yes.
Done.